OpenStudy (anonymous):

Simplify the expression -6+i/-5+i A. 31+i/26 B. 29+i/26 C. 31+i/24 D. 31+11i/26

OpenStudy (anonymous):

@XxNina26xX can you help me or are you just looking around

OpenStudy (anonymous):

I'm trying, give me a a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Rationalise it i.e., multiply the numerator and denominator by -5-i

OpenStudy (anonymous):

ok then what will i came up with my answer

OpenStudy (anonymous):

@Luigi0210 @Agent47

OpenStudy (agent47):

-6+i/-5+i = (-6+i)(-5-i)/(25+1) = (30 + 6i-5i+1)/26 = (31+I)/26

OpenStudy (anonymous):

thank @Agent47